Circular No. 6067

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET MUELLER (1993a) Extension to ephemeris on IAUC 6007, from elements on MPC 23649:

1994 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Aug. 26 20^{h}21.^{m}26 −52^{o}37.^{′}8 2.503 3.264 131.^{o}8 13.^{o}4 13.1
Sept. 5 19 56.97 −53 28.2 2.714 3.350 121.3 14.9 13.4
15 19 38.43 −53 41.3 2.947 3.437 110.9 15.9 13.7
25 19 25.55 −53 33.3 3.192 3.524 101.0 16.2 14.0
Oct. 5 19 17.65 −53 15.2 3.443 3.611  91.5 16.1 14.3
15 19 13.86 −52 53.9 3.694 3.698  82.5 15.5 14.5
25 19 13.35 −52 33.4 3.939 3.786  73.9 14.6 14.8
Nov. 4 19 15.42 −52 15.7 4.174 3.873  65.7 13.5 15.0
14 19 19.54 −52 01.9 4.395 3.961  58.0 12.2 15.2
24 19 25.22 −51 52.4 4.599 4.048  50.8 10.9 15.4
Dec. 4 19 32.11 −51 47.6 4.784 4.135  44.2  9.6 15.6
14 19 39.93 −51 47.7 4.947 4.223  38.6  8.4 15.7
24 19 48.42 −51 53.0 5.087 4.310  34.3  7.4 15.9

Total magnitude estimates (cf. IAUC 5993): June 9.43 UT, 11.7 (A. Hale, near Orogrande, NM, 0.41-m reflector, visual); July 3.72, 12.3 (A. Nakamura, Kuma, Shikoku, Japan, 0.60-m reflector + unfiltered CCD); 15.72, 12.0 (Nakamura).

PERIODIC COMET MACHHOLZ 2 (1994o) The primary component of this comet has evidently brightened since

discovery. P. Pravec reports that 120-s CCD images taken through a 2′ aperture by M. Wolf and L. Šarounová with the Ondřejov 0.65-m reflector (+ V filter) shows that the total brightness was ∼ 2–3 mag brighter on Aug. 23.07 UT than on Aug. 15.92 and 16.90.

Total visual magnitude estimates (cf. IAUC 6057): Aug. 19.07 UT, 8.2 (L. Szentaskó, Veresegyház, Hungary, 0.33-m reflector); 23.09, 7.8 (J. Lancashire, Cambridge, England, 10×80 binoculars); 28.12, 8.6 (H. Mikuž, Ljubljana, Slovenia, 0.20-m f/2 Baker-Schmidt camera + V filter + ST-6 CCD); 28.36, 7.7 (J. E. Bortle, Stormville, NY, 20×80 binoculars); 30.36, 7.6 (Bortle).

1994 August 30 Daniel W. E. Green